# Benchmarking Federated Learning for Semantic Datasets: Federated Scene Graph Generation

### v1

@@ IV-B Data Partition with Semantic Heterogeneity
Dirichlet distribution-based partition: From the strategy suggested in [21], the amount of data each client takes from cluster $k$ is governed by the sampling from the Dirichlet distribution. We can design the non-IID data partition into $U$ clients by sampling multinomial $\mathbf{p}_{u}\sim\mathrm{Dir}_{n}(\bm{\alpha})$ of $u$-th client from Dirichlet distribution with $\bm{\alpha}$, where $\sum^{n}_{i=1}\mathbf{p}_{u,i}=1$ and $\mathbf{p}_{u,i}\in[0,1]$ for all client $u\in\{1,...,U\}$ and for all cluster $i\in\{1,...,n\}$, and $\bm{\alpha}=(\alpha_{1},\alpha_{2},...,\alpha_{n})$ for all $\alpha_{i}\in(0,\infty]$ is a concentration parameter vector. Same as [21], we set all the element of $\bm{\alpha}$ is same, $\alpha_{1}=\alpha_{2}=...=\alpha_{n}$. Each client $u$ can sample training data from a dataset according to a proportion $\mathbf{p}_{u,i}$ without replacing each cluster $i$. The data heterogeneity can be controlled by a value of $\alpha$. As $\alpha$ increases, the homogeneity of the data across clients increases, while as $\alpha$ decreases, the heterogeneity of the data across clients increases.
IV-C Proof-of-concept: FL Benchmark for Panoptic Scene Graph Generation (PSG)
Herein, we provide a proof-of-concept of our FL benchmark process by constructing the FL benchmark for PSG dataset.
@@ IV-C Proof-of-concept: FL Benchmark for Panoptic Scene Graph Generation (PSG)
i) Discovering data clusters: PSG dataset contains object, subject, and predicate labels for each image sample. For simplicity, we utilize 13 object/subject categories and 7 predicate categories, which are the super-classes of fine-grained labels. Due to page limitations, we have uploaded the object/subject and predicate categories we used to our GitHub. Therefore, the dimension of the category tensor is $\mathcal{F}(\mathcal{Y})\in\mathbb{R}^{13\times 13\times 7}$. For the category tensor, we perform $K$-means Clusters to obtain multiple semantic clusters. Consequently, we obtain 5 different clusters with discriminated semantics:
$\mathcal{K}\Big{(}\mathcal{F(Y)}_{1}^{|\mathcal{D}|}\Big{)}\rightarrow\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3},\mathcal{C}_{4},\mathcal{C}_{5}\}.$ (5)
@@
Fig. 4 illustrates the sample images from each cluster and the PCA clustering result image. By examining the samples for each cluster, we observe the following distinguished features for each cluster and the imbalance between clusters:
@@
    This cluster primarily contains images related to animals. We observe that it contains a large number of animal objects compared to others. The predicates are composed of actions that animals trivially perform.
@@
    This cluster is dominated by daily photographs of people, which constitutes the largest portion of PSG dataset. This cluster mainly contains many objects related to daily activities by human beings, such as food photographs that frequently appear in daily life.
@@
    This cluster is composed of sports images. It contains many objects associated with sports and predicates such as ‘playing,’ which are more prevalent than others.
@@
Notably, Clusters 2 and 4 contain somewhat similar images, mainly of ‘people’. However, the predicates in Cluster 4 relate to sports activities, which is clearly discriminated from Cluster 2. Also, Cluster 3 and 5 look similar because of urban landscapes, but Cluster 3 leans to focus on cityscapes with transportation, and Cluster 5 focuses on urban/nature-combined city views. The qualitative visualization clearly demonstrates that our K-means Clustering of Category Tensor effectively and intuitively segments PSG dataset, leading to the splits given semantic information.
ii) Data partition: Based on the discovered 5 semantic clusters, our benchmark provides two options for data distribution: i) Shard-based partitioning and ii) Dirichlet distribution-based partitioning. Regardless of the choice of partitioning, when partitioning becomes heterogeneous, the data distribution at clients strongly deviates in the sense of semantic clusters, which leads to strong semantic heterogeneity. Otherwise, the data distribution of clients becomes homogeneous, yielding evenly distributed semantic information. The data partitioning is quite straightforward, so we will further describe the detailed settings in the following Section V.
V Experiments: Benchmarks for PSG in FL
V-A Experiment Settings
We extensively evaluate the existing panoptic scene graph generation (PSG) models on our benchmark, including the following methods: IMP [39], MOTIFS [40], VCTree [42], and GPS-Net [55]. In these PSG methods, we use the same pretrained object detector Faster R-CNN[64]. The communication cost is the crucial factor in FL scenarios; we freeze the pretrained object detector and focus on predicate classification. Therefore, each client exclusively trains and aggregates the relation head, which is responsible for processing predicates.
Dataset description: For the detailed comparison of experiments, we utilize PSG dataset [47] which includes diversified images with rich relational annotations, where each image is annotated with objects, panoptic segmentation masks, and fine-grained relationships between those objects. It not only identifies individual objects and their relationships but also includes stuff (amorphous background regions like “sky” or “grass”), which is often overlooked in other datasets. PSG dataset leverages VG150 [65] and COCO [66] datasets, which are not perfect but popularly used in scene graph generation tasks, by integrating their comprehensive object and relationship annotations into a more focused dataset designed specifically for the scene graph generation. To be specific, PSG dataset directly inherits the panoptic segmentation annotations from COCO. On the other hand, the VG150 dataset contains many ‘trivial’ (not meaningful) predicates with the direction of predicates (e.g., of in hair-‘of’-man, has in man-‘has’-head), and PSG dataset gets rid of these predicates. PSG dataset contains 133 objects and 374 relationships, sufficiently covering the diversity from the VG150 and COCO datasets while compensating for their limitations. These relationships are more detailed and extensive than in other datasets, allowing for richer scene graph representations.
TABLE I: Dataset Information
  Data Amount
PSG dataset - Train 46 K
  Cluster Balanced (24.5%) Imbalanced (100%)
Cluster 0 2.2K (4.9%) 2.2K (4.9%)
Cluster 1 2.2K (4.9%) 27K (58.1%)
Cluster 2 2.2K (4.9%) 5.1K (10.9%)
Cluster 3 2.2K (4.9%) 3.3K (7.1%)
Cluster 4 2.2K (4.9%) 8.8K (19.0%)
The training dataset comprises 46,563 images. Table I describes the information about the dataset we have composed with Category Tensor K-means Clustering described in Fig. 3. An imbalanced dataset refers to a dataset where the number of data points in each cluster is not equal after clustering. As mentioned in Sec IV-B, to eliminate cluster imbalance, we randomly sampled the data from each cluster to match the quantity of the smallest cluster, Cluster 0. This process ensured that all clusters had the same amount of data (2.2K images), resulting in a balanced dataset.
Baseline description: We chose four panoptic scene graph generation models to validate our benchmark. For ease of understanding, we present the overview of PSG methods in Fig. 5. (a) IMP[39] (Iterative Message Passing) is a method in scene graph generation that repeatedly delivers messages for each node (object) and edge (relationships) and updates information to optimize objects and relationships simultaneously with GRUs. (b) MOTIFS[40] leverages the role of motifs in object relationships to guide scene graph generation, using bi-LSTMs to model context for better relationship prediction. (c) VCTree [42] constructs a dynamic tree structure to hierarchically capture visual and semantic dependencies among objects using the bidirectional TreeLSTM, enabling effective representation of relationships for scene graphs. (d) GPS-Net[55] enhances scene graph generation by integrating the Graph Property Sensing module that dynamically captures both global and local contextual properties of the graph. The Graph Property Sensing module consists of (1) Direction-aware Message Passing (DMP), (2) Node Priority Sensitive Loss (NPS-loss) utilizing object classification scores $p_{i}$ and (3) Adaptive Reasoning Module (ARM). It leverages an adaptive message-passing mechanism to encode object relationships better and address biases in the data, resulting in more accurate and context-aware scene graphs.
Experiment setups: We set up an FL scenario with one server and 100 clients, distributing the training data of the existing PSG dataset [47] to the 100 clients. The test data for our benchmark was the same as PSG test dataset. In each round, five active clients were randomly selected, and the test data was evaluated using the aggregated global model from the server. Each client performs local training with one epoch and a batch size 16. The total number of training rounds reaches up to 100, and we report the R/mR@K performance of the final averaged model. Following the benchmark in [47], we set the SGD optimizer to local optimizer with a learning rate of 0.02, momentum of 0.9, weight decay of 0.0001, and gradient clipping with a max L2 norm of 35.
TABLE II: Comparison of the performances of PSG methods on the proposed FL benchmark
@@ V-A Experiment Settings
Benchmark setups: As described in Subsection IV-B, to ease the cluster imbalance, we randomly sampled data from each cluster to ensure an equal amount of data for each cluster. We tested 6 types of data partitioning as follows:
(1) Random, where data is distributed randomly among all clients, ensuring nearly equal sizes for each.
(2) Shard-based partition IID, where we set $p=5$, where $p$ is the number of clusters that client sample from. As aforementioned, when $p$ equals the number of clusters, the data from each cluster is equally distributed among 100 clients.
(3) Shard-based partition non-IID, where we set $p=1$ for imposing semantic heterogeneity. Each cluster is assigned 20 clients, and all clients have the same amount of data.
(4), (5) and (6) Dirichlet distribution-based partition ranging from an IID case to a strong non-IID case, where we tested three different levels of semantic heterogeneity by using $\alpha=[10,1,0.2]$.
Metrics: By following the work of [47] that has first suggested the PSG task, we use ‘Recall@K (R@K)’ and ‘mean Recall@K (mR@K)’ as the performance metrics, which respectively calculate the triplet recall and mean recall for every predicate category, given the top K triplets from a PSG method. $K$ varies from 20 to 100. Moreover, R@K is dominated by high-frequency relations, and mR@K assigns equal weight to all relation classes. In datasets with severe long-tailed problems, e.g., PSG dataset, mR@K can provide more meaningful insights into model performance.
V-B In-depth Analysis
@@
Results: Table II shows the test accuracy on the test set of PSG dataset. We have focused on the Mean Recall (‘mR’) performance. Also, we focus on the most challenging case with $K=20$.
@@
i) CL vs. IID. The performance has been mostly degraded when comparing CL and IID cases. The averaged gaps for mR@20 are $-2.45\%$ and $-2.71\%$, for ‘Shard-IID’ and ‘Dir($\alpha=10$)’, respectively. Each client has approximately 114 images, and due to the limited data, there appears to be a performance difference between the CL and IID scenarios. CL can collectively form a mini-batch across clients, but IID forms a mini-batch per client in a decentralized manner.
@@
ii) IID vs. Random. When data is randomly divided, it will tend to have a distribution close to IID so that there is a minimal performance drop. The averaged gaps for mR@20 are $-0.32\%$ and $-0.12\%$, for ‘Shard-IID’ and ‘Dir($\alpha=10$)’, respectively. The results confirm that the random partitioning naively conducted in prior studies is unsuitable for imposing semantic heterogeneity, showing similar results as the IID case.
@@
iii) IID vs. non-ID. We confirm large performance degradations in most cases. First, in the case of a shard-based partition, the averaged gap for mR@20 is $-0.77\%$. Second, in the case of the Dirichlet distribution-based partition, i.e., comparing Dir($\alpha=10$) and Dir($\alpha=0.2$), the averaged gap for mR@20 is shown to be $-0.64\%$. The performance drops from IID to non-IID reveal that PSG methods struggle to aggregate a global model when facing a string semantic heterogeneity. Also, when a PSG method shows a minimal performance drop due to the heterogeneity, it directly demonstrates the robustness against semantic heterogeneity. MOTIFS shows the outliers in mR, where the moderate non-IID case ($\alpha=1$) compared to the non-IID case ($\alpha=0.2$) shows minimal differences: 4.09% vs. 4.28% in mR@20, and 4.62% vs. 4.71% in mR@100. Although the results may seem unexpected, the differences are not significant. Notably, when $\alpha=10$, corresponding to the IID case, shows the best performance: 4.39% in mR@20 and 4.86% in mR@100, aligning with our expectations. Based on this observation, we conjecture that the behavior at the moderate non-IID can be a little shaky in a few cases, but it behaves as expected in the IID case.
@@
PSG Model comparisons: We here to discuss the robustness of the existing PSG methods against semantic heterogeneity. We conclude that IMP is shown to be relatively vulnerable in handling semantic heterogeneity in FL, i.e., a large gap of $-1.58\%$ for mR@20 is observed when comparing Dir($\alpha=10$) and Dir($\alpha=0.2$). Compared to others, it has a relatively smaller model architecture and suffers from the long-tailed problem in PSG dataset. We conjecture that the aspects of IMP lead to considerable performance drops in our non-IID testing. VCTree includes a tree construction process trained through reinforcement learning, resulting in a more complex model structure compared to MOTIFS. Consequently, in a FL scenario with small-scale client data, the performance of VCTree is degraded. GPS-Net employs key elements, e.g., DMP, NPS-loss, ARM, to resolve the long-tailed problem. We conjecture that it yields the outperforming results of GPS-Net in our FL benchmarks. In FL, clients have a small number of data samples, which makes worse the long-tailed problem. Because GPS-Net has two key factors that pay more attention to objects and predicates with smaller occurrences, it shows remarkable performances.
V-C Convergence Behavior

### v2

@@ 3.2 Data Partition with Semantic Heterogeneity
Dirichlet distribution-based partition: From the strategy suggested in [1], the amount of data each client takes from cluster $k$ is governed by the sampling from the Dirichlet distribution. We design the non-IID data partition into $U$ clients by sampling a multinomial probability vector for each client $u$, denoted as: $\mathbf{p}_{u}\sim\mathrm{Dir}_{n}(\boldsymbol{\alpha})$ where the probability vector is $\mathbf{p}_{u}=(p_{u,1},p_{u,2},\ldots,p_{u,n})$, and the concentration parameter of Dirichlet distribution is $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})$, for $\alpha_{i}>0,\forall i\in\{1,2,\ldots,n\}$. Each $u$th client samples training data from a dataset according to the proportion $p_{u,i}$ without replacement for each cluster $i$. The data heterogeneity is controlled by $\boldsymbol{\alpha}$: As $\boldsymbol{\alpha}\to\infty$, the data distribution goes to IID; in contrast, as $\alpha\to 0$, it goes to non-IID.
3.3 Proof-of-concept: FL Benchmark for Panoptic Scene Graph Generation (PSG)
We provide a proof-of-concept of our FL benchmark process by constructing the FL benchmark for PSG dataset.
@@ 3.3 Proof-of-concept: FL Benchmark for Panoptic Scene Graph Generation (PSG)
(i) Discovering data clusters: PSG dataset contains object, subject, and predicate labels for each image sample. For simplicity, we utilize 13 object/subject categories and 7 predicate categories, which are the super-classes of fine-grained labels Therefore, the dimension of the category tensor is $\mathcal{F}(\mathcal{Y})\in\mathbb{R}^{13\times 13\times 7}$. We perform $K$-means Clustering for the category tensor to obtain multiple semantic clusters, obtaining 5 clusters with discriminated semantics. [^1] We attach the detailed description in Appendix A.:
$\mathcal{K}\Bigl{(}\mathcal{F(Y)}_{1}^{|\mathcal{D}|}\Bigr{)}\rightarrow\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3},\mathcal{C}_{4},\mathcal{C}_{5}\}.$ (2)
(ii) Data partition: Based on the discovered 5 semantic clusters, our benchmark provides two options for data distribution: (i) Shard-based partitioning and (ii) Dirichlet distribution-based partitioning. As partitioning becomes heterogeneous, the data distribution at clients strongly deviates in the sense of semantic clusters. Otherwise, the data distribution of clients becomes homogeneous, yielding evenly distributed semantic information.
4 Experiments: Benchmarks for PSG in FL
4.1 Experiment Settings
We evaluate the existing panoptic scene graph generation (PSG) models on our benchmark with the following methods: IMP [27], MOTIFS [29], VCTree [26], and GPS-Net [20]. Because these PSG methods use the same pretrained object detector Faster R-CNN [7], and the communication cost is crucial in FL scenarios, we freeze the pretrained object detector and focus on predicate classification. Therefore, each client trains and aggregates the relation head, responsible for processing predicates by capturing the semantic relationships between objects.
An imbalanced dataset contains a number of data points in each cluster that is not equal after clustering. We randomly sampled the data from each cluster to match the quantity of the smallest cluster to eliminate cluster imbalance. This process ensured that all clusters had the same data (2.2K images), resulting in a balanced dataset.
Experiment setups: We set up an FL scenario with one server and 100 clients, distributing the training data of the existing PSG dataset [28] to the 100 clients. The test data for our benchmark is the same as the PSG test dataset. Five active clients are randomly selected in each round, and the test data is evaluated using the aggregated global model from the server. Each client performs local training with one epoch and a batch size 16. The total number of training rounds is 100, and we report the R/mR@K performance of the final averaged model. Following the benchmark in [28], we set the SGD optimizer to a local optimizer with a learning rate of 0.02, momentum of 0.9, weight decay of 0.0001, and gradient clipping with a max L2 norm of 35.
@@ 4.1 Experiment Settings
Clustered PSG Dataset Description: By examining the samples for each cluster, we observe the following features for each cluster and the imbalance between clusters [^2] The visualization of the clustering is in Appendix A.2:
@@
    We observe that it contains a large number of animal objects compared to others. The predicates are composed of actions that animals trivially perform.
@@
    This cluster is dominated by daily photographs of people, which constitutes the largest portion of PSG dataset. This cluster is mainly related to daily activities by human beings that frequently appear in daily life.
@@
    This cluster is composed of sports and kinetic images, containing predicates such as ‘playing,’ which are more prevalent than others.
@@
Notably, Clusters 2 and 4 contain somewhat similar images, mainly of ‘people’. However, the predicates in Cluster 4 relate to sports, clearly distinguishing it from Cluster 2. Also, Cluster 3 and 5 look similar because of urban landscapes, but Cluster 3 tends to focus on cityscapes with transportation, and Cluster 5 focuses on urban/nature-combined views.
Table 1: Comparison of the performances of PSG methods on the proposed FL benchmark
@@
Benchmark setups: We randomly sampled data from each cluster to ensure an equal amount of data for each cluster to ease the cluster imbalance. We test 6 types of data partitioning as follows:
(2) Shard-based partition IID: We set $p=5$, where $p$ is the number of clusters that client sample from. When $p$ equals the number of clusters, the data from each cluster is equally distributed among 100 clients.
(3) Shard-based partition non-IID: We set $p=1$ for imposing semantic heterogeneity. Each cluster is assigned 20 clients, and all clients have the same amount of data.
(4), (5) and (6) Dirichlet distribution-based partition: We test three levels of semantic heterogeneity by using $\alpha=[10,1,0.2]$ to simulate from an IID to a non-IID case.
Metrics: By following [28] which suggested the PSG task, we use ‘Recall@K (R@K)’ and ‘mean Recall@K (mR@K)’ as the performance metrics, which calculate the triplet recall and mean recall for every predicate category, given the top $K\in[20,100]$ triplets from a PSG method. Moreover, R@K is dominated by high-frequency relations, and mR@K assigns equal weight to all relation classes. In datasets with severe long-tailed problems, e.g., PSG dataset, mR@K can provide more meaningful insights into model performance.
4.2 In-depth Analysis
@@
[…]
@@
(i) CL vs. IID. The performance has been mostly degraded when comparing CL and IID cases. The averaged gaps for mR@20 are $2.45\%$ and $2.71\%$, for ‘Shard-IID’ and ‘Dir($\alpha=10$)’. Each client has approximately 114 images, and due to the limited data, there appears to be a performance difference between the CL and IID scenarios. CL can collectively form a mini-batch across clients, but IID forms a mini-batch per client in a decentralized manner.
@@
(ii) IID vs. Random. When data is randomly divided, it will tend to have a distribution close to IID so that there is a minimal performance drop. The averaged gaps for mR@20 are $0.32\%$ and $0.12\%$, for ‘Shard-IID’ and ‘Dir($\alpha=10$)’, respectively. The results confirm that the random partitioning naively conducted in prior studies is unsuitable for imposing semantic heterogeneity, showing similar results as the IID case.
@@
(iii) IID vs. non-ID. We confirm large performance degradations in most cases. First, in the case of a shard-based partition, the averaged gap for mR@20 is $0.77\%$. Second, in the case of the Dirichlet distribution-based partition, i.e., comparing Dir($\alpha=10$) and Dir($\alpha=0.2$), the averaged gap for mR@20 is shown to be $0.64\%$. The performance drops from IID to non-IID reveal that PSG methods struggle to aggregate a global model under strong semantic heterogeneity. MOTIFS shows the outliers in mR, where the moderate non-IID case ($\alpha=1$) compared to the non-IID case ($\alpha=0.2$) shows minimal differences: 4.09% vs. 4.28% in mR@20, and 4.62% vs. 4.71% in mR@100. It looks unexpected, but it is not a considerable amount. Also, we want to point out that when $\alpha=10$, which is the IID case, the performance becomes maximized: 4.39% in mR@20 and 4.86% in mR@100, which coincides with our expectations. We conjecture that the behavior at the moderate non-IID can be a little shaky in a few cases, but it finally behaves as expected in the IID case. Although the results may seem unexpected, the differences are not significant. Notably, when $\alpha=10$, corresponding to the IID case, shows the best performance: 4.39% in mR@20 and 4.86% in mR@100, aligning with our expectations. Based on this observation, we conjecture that the behavior at the moderate non-IID can be a little shaky in a few cases, but it behaves as expected in the IID case.
@@
PSG Model comparisons: We discuss the robustness of the existing PSG methods against semantic heterogeneity. We conclude that IMP is relatively vulnerable in handling semantic heterogeneity in FL, i.e., a large gap of $1.58\%$ for mR@20 is observed when comparing Dir($\alpha=10$) and Dir($\alpha=0.2$). It has a smaller model architecture and suffers from the long-tailed problem in the PSG dataset. We conjecture that the aspects of IMP lead to notable performance drops in our non-IID testing. VCTree includes a tree construction process trained via reinforcement learning, resulting in a more complex model structure than MOTIFS. Consequently, in the FL scenario with small-scale client data, its performance is degraded. Because GPS-Net employs key elements, e.g., DMP, NPS-loss, and ARM, to resolve the long-tailed problem, we conjecture that it yields the outperforming results of GPS-Net in our FL benchmarks.
4.3 Extension to FL algorithms
Table 2: Comparison of the FedAvg, FedAvgM, FedAdam performances of PSG methods.
  R/mR@K Method FedAvg FedAvgM[8] FedAdam[24]
Shard non-IID Shard non-IID Shard non-IID
R/mR@20 IMP 11.26 / 2.28 13.23 / 3.83 (+1.55%) 13.32 / 4.78 (+2.50%)
MOTIFS 13.33 / 4.06 15.47 / 5.80 (+1.74%) 15.89 / 5.56 (+1.50%)
VCTree 12.49 / 3.99 15.39 / 5.66 (+1.67%) 15.53 / 5.09 (+1.10%)
GPS-Net 14.57 / 5.90 16.18 / 5.91 (+0.01%) 15.66 / 5.98 (+0.08%)
R/mR@50 IMP 12.57 / 2.59 14.73 / 4.24 (+1.65%) 15.03 / 5.41 (+2.82%)
MOTIFS 14.92 / 4.48 17.23 / 6.23 (+1.75%) 17.66 / 6.01 (+1.53%)
VCTree 14.04 / 4.41 17.02 / 6.10 (+1.69%) 16.95 / 5.40 (+0.99%)
GPS-Net 16.37 / 6.36 18.00 / 6.33 (-0.03%) 17.31 / 6.42 (+0.06%)
R/mR@100 IMP 13.06 / 2.68 15.32 / 4.38 (+1.70%) 15.63 / 5.57 (+2.89%)
MOTIFS 15.43 / 4.65 17.83 / 6.38 (+1.73%) 18.21 / 6.14 (+1.49%)
VCTree 14.62 / 4.54 17.58 / 6.30 (+1.76%) 17.42 / 5.53 (+0.99%)
GPS-Net 16.91 / 6.49 18.68 / 6.52 (-0.03%) 17.90 / 6.55 (+0.06%)
$(\cdot)$ indicates the difference in mR@K when each algorithm is applied compared to FedAvg.
Next, we verify whether the improvements in FL algorithms stay valid in our benchmark. We conducted additional experiments on two FL algorithms employing momentum. Momentum-based update strategies prove effective in maintaining local training closer to the global update direction. By mitigating the adverse effects of data distribution discrepancies, these algorithms can enhance both convergence stability and model performance.
4.3.1 FedAvgM
[…]
$w_{g}^{r+1}$ $=w_{g}^{r}-v^{r},$ (3)
$v^{r}$ $=\beta v^{r-1}+\sum_{k=1}^{K}\frac{n_{k}}{n}\Delta w_{k}^{r}$ (4)
Results: FedAvgM sufficiently improves the performance of all methods. For R/mR@20, R/mR@50, and R/mR@100, there are average performance improvements of +1.24%, +1.27%, and +1.30% in the Shard-nonIID case, respectively. These performance improvements under FedAvgM can be attributed to FedAvgM’s momentum-based updates which help stabilize training by reducing local update oscillations in the optimization process, leading to faster convergence and better generalization across heterogeneous client data. IMP, MOTIFS, and VCTree showed noticeable performance increases, while GPS-Net did not. GPS-Net in the Shard non-IID case shows a negligible gap (i.e., $\leq 0.03\%$), indicating that GPS-Net already incorporates factors that mitigate the effects of heterogeneity.
4.3.2 FedAdam
FedOpt [24] provides a framework for improving optimization, which supports server-side optimization algorithms (e.g., Adam) to enhance convergence and stability. This approach deals with divergent client data distributions and fluctuations in client participation rates. FedOpt uses different optimizers in local and global updates. In our case, we utilize the Adam [10] optimizer for global updates:
$m^{r+1}$ $=\beta_{1}m^{r}+(1-\beta_{1})\Delta w^{r},$ (5)
$v^{r+1}$ $=\beta_{2}v^{r}+(1-\beta_{2})(\Delta w^{r})^{2},$ (6)
$w_{g}^{r+1}$ $=w_{g}^{r}-\eta\frac{m^{r+1}}{\sqrt{v^{r+1}}+\epsilon},$ (7)
[…]
Results: FedAdam demonstrated a marginally superior performance improvement compared to FedAvgM. For R/mR@20, R/mR@50, and R/mR@100, there are average performance improvements of +1.30%, +1.35%, and +1.36% in the Shard-nonIID case, respectively. Interestingly, FedAdam shows a noticeable performance improvement when combined with IMP, as shown in the table. IMP has a lower initial performance (based on R/mR@20, R/mR@50, and R/mR@100) when compared to other methods (MOTIFS, VCTree, GPS-Net). However, when combined with FedAdam, it showed the greatest performance improvement (+2.89% in R/mR@100 case). IMP simply learns by iteratively updating relationships between objects. IMP is prone to learning by being overly head-class-biased in class imbalances, and performance degradation is inevitable in tail classes. In this environment, FedAdam has most likely improved its model effectively in the tail class, where losses are concentrated due to the long-tailed problem. Contrary to IMP, GPS-Net has various strategies to solve the long-tailed problem. Similar to the FedAvgM experimental result, GPS-Net showed no significant change in performance.
We can conclude that the enhancement of FL algorithms is effective when dealing with scenarios involving diverse semantic information across clients. Furthermore, the experimental outcomes with IMP and GPS-Net reveal an intriguing connection: the long-tailed problem encountered in scene graph generation tasks shares notable similarities with the data heterogeneity issues faced in FL. In scenarios where scene graph generation tasks must be addressed in a distributed data environment, selecting a scene graph generation method with a strong strategy for handling long-tailed problems or choosing an FL algorithm that effectively deals with data heterogeneity would significantly increase the likelihood of simultaneously tackling both challenges.
Table 3: Comparisons of FL algorithms on CelebA emotion classification (Smiling vs. Not smiling) task
  Method Shard Dirichlet distribution
IID non-IID $\alpha=10$ $\alpha=1$ $\alpha=0.2$
FedAvg 90.94 (32) 90.43 (52) 91.56 (32) 91.65 (37) 91.03 (47)
FedAvgM 92.93 (27) 91.22 (32) 93.4 (27) 92.93 (27) 91.32 (37)
FedAdam 91.84 (52) 91.03 (52) 91.71 (52) 91.77 (52) 90.79 (52)
($\cdot$) is the communication rounds to reach 85% Acc.
5 Additional Experiments on CelebA
To demonstrate the generalizability, we also applied our clustering method to the CelebA dataset, which comprises 40 attributes (e.g., Eyeglasses, Wearing hat, Wavy hair), not a specific label. [^4] We attach the details on CelebA experiments in the Appendix. E
5.1 Construction of non-IID CelebA Dataset
In prior works, the CelebA dataset is distributed only according to identity in a non-IID case, e.g., each client has pictures of the same person, which limits the method for representing a kind of non-IID setting. It also cannot represent practical cases, such as the contents of CCTV or broadcast systems. Therefore, we build a non-IID dataset with the following steps. First, we perform K-Means clustering with 5 clusters. Since the attributes consist of binary values (-1, 1), we do not apply super-classes to preserve the meaning of each attribute. Second, we cluster images and allocate the cluster index to each image. Based on the cluster index of each image, we build non-IID cases by applying data partitioning methods such as shard-based and Dirichlet distribution-based partitions. We evaluate the classification task to check whether each person is smiling in 3.
5.2 Analysis
[…]
6 Conclusion
Our work takes a decisive step toward closing the gap between federated learning (FL) research and high-level visual understanding by introducing the first benchmark framework for multi-semantic vision tasks under controlled data heterogeneity. Existing approaches for generating data heterogeneity rely on single-label datasets and cannot be extended to multiple semantic collections. To overcome this limitation, we propose a clustering-based scheme that groups samples based on the semantics of the images. This design naturally accommodates datasets with multiple labels per image, such as PSG and CelebA. We validate our framework on the PSG dataset and CelebA by partitioning the data according to cluster assignments and measuring convergence behavior. As the data heterogeneity increases, i.e., from non-IID to IID, training converges more slowly and shows lower final accuracy for both datasets, confirming the ability to construct heterogeneity. Furthermore, our benchmark demonstrated consistent trends with prior FL studies, even when extended to various FL scenarios, i.e, changes in participation rates and number of clusters.
While the proposed approach shows significant results, it has limitations. First, determining the optimal number of clusters remains challenging. Second, the method assumes the availability of semantic annotations, which may not always be accessible.
Nevertheless, the proposed benchmark can be adopted in practical use. It can be extended to multi-label and structured data tasks such as multimodal learning, visual question answering, and relation extraction. The framework is highly applicable to domains such as broadcasting and media, where centralized learning is often impractical due to stringent requirements regarding raw data confidentiality, intellectual property, and content ownership. FL with semantic-aware partitioning offers a collaborative model training without compromising sensitive media assets.
Acknowledgments
This work was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2021-0-00852, Development of Intelligent Media Attributes Extraction and Sharing Technology) and National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (No. RS-2024-00459023).
[…]
Appendix A Detailed Desciption of the Dataset
A.1 PSG dataset
We utilize the PSG dataset [28] where each image is annotated with objects (including stuff such as “sky" or “grass"), panoptic segmentation masks, and fine-grained relationships between those objects. PSG dataset leverages VG150 [12] and COCO [19] datasets by integrating their comprehensive object and relationship annotations into the scene graph generation task. Specifically, the PSG dataset directly inherits the panoptic segmentation annotations from COCO. The VG150 dataset contains many ‘trivial’ (not meaningful) predicates with the direction of predicates (e.g., of in hair-‘of’-man, has in man-‘has’-head), and the PSG dataset gets rid of these predicates. The PSG dataset contains 133 objects and 374 relationships, sufficiently covering the diversity from the VG150 and COCO datasets while compensating for their limitations. These relationships are more detailed and extensive than in other datasets, allowing for richer scene graph representations.
Table 4: Dataset Information
  Data Amount
PSG dataset - Train 46 K
  Cluster Balanced (24.5%) Imbalanced (100%)
Cluster 0 2.2K (4.9%) 2.2K (4.9%)
Cluster 1 2.2K (4.9%) 27K (58.1%)
Cluster 2 2.2K (4.9%) 5.1K (10.9%)
Cluster 3 2.2K (4.9%) 3.3K (7.1%)
Cluster 4 2.2K (4.9%) 8.8K (19.0%)
A.2 Clustered PSG Dataset
Fig. 3 illustrates the sample images from each cluster and the PCA clustering result image. The qualitative visualization clearly demonstrates that our K-means Clustering of Category Tensor effectively and intuitively segments PSG dataset, leading to the splits given semantic information.
Appendix B Description of Baseline Algorithms
We choose four panoptic scene graph generation models to validate our benchmark. For ease of understanding, we present the overview of PSG methods in Fig. 4. (a) IMP[27] (Iterative Message Passing) is a method in scene graph generation that repeatedly delivers messages for each node (object) and edge (relationships) and updates information to optimize objects and relationships simultaneously with GRUs. (b) MOTIFS[29] leverages the role of motifs in object relationships to guide scene graph generation, using bi-LSTMs to model context for better relationship prediction. (c) VCTree [26] constructs a dynamic tree structure to hierarchically capture visual and semantic dependencies among objects using the bidirectional TreeLSTM, enabling effective representation of relationships for scene graphs. (d) GPS-Net[20] enhances scene graph generation by integrating the Graph Property Sensing module that dynamically captures both global and local contextual properties of the graph. The Graph Property Sensing module consists of (1) Direction-aware Message Passing (DMP), (2) Node Priority Sensitive Loss (NPS-loss) utilizing object classification scores $p_{i}$ and (3) Adaptive Reasoning Module (ARM). It leverages an adaptive message-passing mechanism to encode object relationships better and address biases in the data, resulting in more accurate and context-aware scene graphs.
Appendix C Additional Experiments
C.1 Convergence Behavior